OpenStudy (anonymous):

I had this problem, (3x^-1/4y^-1)^-2 and i get a fraction within a fraction within a fraction and i am having trouble simplifying the ^ means raised to the power of,,, please help

OpenStudy (anonymous):

first invert it because its a negative power then square both denomniator and numerator

myininaya (myininaya):

\[(\frac{4y^{-1}}{3x^{-1}})^2\] that 2 is suppose to be an exponent this should help

OpenStudy (anonymous):

so its the same as 1/((3y/4x)^2)

myininaya (myininaya):

\[(\frac{4x}{3y})^2\]

myininaya (myininaya):

\[\frac{16x^2}{9y^2}\]

OpenStudy (anonymous):

(16 x^2)/(9 y^2)

OpenStudy (anonymous):

(9y^2)/(16x^2) is the correct answer because in the original problem its all to the power of -2 which means its flipped

myininaya (myininaya):

rocket ur answer is wrong unless he wrote it wrong

OpenStudy (anonymous):

x^-1 is the same as 1/x

myininaya (myininaya):

rockets answer would be right if we had ((3y)^(-1))/((4x)^(-1)))^(-2)

OpenStudy (anonymous):

i thought that is what we had myininaya

OpenStudy (anonymous):

\[(4y ^{-1}\div3x ^{-1})^{2} \]

OpenStudy (anonymous):

rocket science is correct it is just based on rule of indices

OpenStudy (anonymous):

how do you remember the negitives from that?

OpenStudy (anonymous):

ok i was looking in the original post, I had this problem, (3x^-1/4y^-1)^-2 and i get a fraction within a fraction within a fraction and i am having trouble simplifying the ^ means raised to the power of,,, please help.

myininaya (myininaya):

ok maybe this would look prettier to you\[(\frac{3x^{-1}}{4y^{-1}})^{-2}=\frac{3^{-2}x^{2}}{4^{-2}y^{2}}=\frac{16x^2}{9y^2}\]

myininaya (myininaya):

is this better looking?

OpenStudy (anonymous):

ok i was looking in the original post, I had this problem, (3x^-1/4y^-1)^-2 and i get a fraction within a fraction within a fraction and i am having trouble simplifying the ^ means raised to the power of,,, please help.

OpenStudy (anonymous):

thank you guys, I get it now

myininaya (myininaya):

i will put one more step in \[\frac{4^2x^2}{3^2y^2}=\frac{16x^2}{9y^2}\]

OpenStudy (anonymous):

thank you